OpenStudy (anonymous):

Determine if 3 * 41/6 is the simpliest form of 29161/6 . If not, tell why not and write the expression in simplest form.

OpenStudy (anonymous):

Definitely not. Because 3 * 41 = 123 and that is nowhere close to 29161 (assuming this isn't a typo). Also 29161/6 is irreducible.

OpenStudy (anonymous):

ITS IS A FRACTION 3 MULTIPLIES BY 4 AND 1 SIXTH

OpenStudy (anonymous):

\(3*4\frac{1}{6}\)?

OpenStudy (anonymous):

YES AND THE OTHER IS THE SAME 2916 1/6

OpenStudy (anonymous):

Still no. \(4\frac{1}{6}\) is \(\frac{25}{6}\) * 3 = \(75/6\). 2916 1/6 is much, much larger. If it is \[2916\frac{1}{6}\] it equals 486 though.

OpenStudy (anonymous):

THANK YOU